If the density of seawater is 1.03 g/cm3

at the surface, what is its density at a depth where

the pressure is 109 dynes/cm2?


1
Expert's answer
2021-03-21T11:33:46-0400

p=109 dynes/cm^2=10.9 Pa,

"h=\frac{p}{\rho g}=\frac{10.9}{1030\cdot 9 .8}=1.08 ~mm,"

This depth is shallow, so destinies are equals.